OpenStudy (chrismatias):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find f(-1) + 5g(-3).

sam (.sam.):

If \[f(x) = x^2 + 1\] \[f(-1) = (-1)^2 + 1\] ---------------------- \[g(x) = 3x + 1\] \[g(-3) = 3(-3) + 1\] ================ Then \[f(-1)+5g(-3)\] \[=((-1)^2+1)+5(3(-3)+1)\]

OpenStudy (whitakerjasmine09):

i dont get it

sam (.sam.):

If f(x) = x^2 + 1, when x=-1, the function f(-1) will be equal to \((-1)^2+1\) Same for g(x), If g(x) = 3x+1, when x=-3, the function g(-3) will be equal to \(3(-3)+1\)

sam (.sam.):

You've been asked to find f(-1) + 5g(-3). So, we knew f(-1) is \((-1)^2+1\) and g(-3) is \(3(-3)+1\) \[f(-1)+5g(-3) \\ \\ =(-1)^2+1+5(3(-3)+1) \\ \\ =?\]